OpenStudy (arbyj):

please help me-- this question's a bit lengthy so ill say the rest in the comments, its concerning grading systems

OpenStudy (arbyj):

test grades: 70, 87, 97 hw grades: 90, 85, 70, 94, 87, 78 particip. grades: 94, 95, 100 99 test avg: 50% homework: 20% par.: 10% final exam: 20% its asking to calculate the lowest grade i can score on my final to get a 90 for the course, how do i calculate it?

OpenStudy (vuriffy):

If only I was a teacher and had their program and just type all this in real quick. Haha. >.<

OpenStudy (vuriffy):

I have to get off soon, but it says test avg: so I would assume you would want to calculate that value to prepare to work out this question fully. 70 + 87 + 97 = 254/3(amount of tests) = 84.66[repeated](test avg)

OpenStudy (arbyj):

no i already have it, those percentages make up my final grade

OpenStudy (arbyj):

my teacher has us set it up like this:

OpenStudy (arbyj):